# Can stellar physics explain GW231123?

### v1

@@ Introduction
The bottom panel of Fig. 1 presents these same models in the BH spin–mass plane. Stars with a helium core mass of $160{\rm M}_{\odot}$ cannot form BHs with a spin above a certain threshold, which depends on the $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ reaction rate. In particular, BHs with spins and masses compatible with the $90\%$ confidence interval of the primary in GW231123 cannot be formed if the $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ reaction rate is smaller than its measured median value.
@@
As anticipated above, our results can be explained by a delay in the onset of photodisintegration instability for fast rotation. As shown in Fig. 2, rotating stars reach lower core temperatures $T_{c}$ before the onset of explosive oxygen burning leads to rapid expansion and a temperature drop. Thus, these stars avoid reaching the conditions required for photodisintegration-induced collapse, and instead undergo PISN. For the median $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate, we find that this occurs at $\Omega_{\rm PISN}\sim 0.4\Omega_{\rm crit}$.
@@
Figure 2: Central temperature $T_{c}$ versus central density $\rho_{c}$ for the simulations described in text, with $M_{\rm in}=160{\rm M}_{\odot}$ for the median $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate. Non-rotating models reach higher core densities and temperatures, crossing into the regime where photodisintegration reactions lead to gravitational collapse, here assumed to be $T_{c}=9\times 10^{9}\rm\,K$. Rotating models, by contrast, terminate at lower $T_{c}$ and $\rho_{c}$, thereby avoiding collapse and instead undergoing a PISN.
Implications for GW231123
@@
Our simulations imply that it is plausible to interpret GW231123’s primary component ($m_{1}\simeq 137{\rm M}_{\odot}$) as having formed via photodisintegration instability collapse. A similar claim was previously made [41] for the primary BH in GW190521, the former high‐mass record holder among GW events, although that interpretation depends strongly on the adopted prior distribution when analyzing the underlying GW data.
@@
In our models, the $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ reaction rate needed to form GW231123’s primary is both broad and realistic, ranging from the median to rates $3\sigma$ stronger. The upper end is compatible with previous hints that higher rates are needed to explain GW observations e.g., the $35{\rm M}_{\odot}$ peak in the BH mass function [28].
Ostensibly, the combination of large masses and large spins in GW231123 could be explained by binaries formed through hierarchical mergers (for context, see Ref. [42] and references therein) and a few studies have already explored this scenario [1, 2, 4]. We argue that it is somewhat unlikely. Hierarchical mergers in dense stellar environments produce a spin distribution that is sharply peaked at $\chi\sim 0.7$ [43, 44], which matches the BH spin in GW231123 only near the lower edge of its 90% credible interval. Spins of $\chi_{1}\sim 0.9$ are too large to be explained as merger remnants [45], requiring substantial fine‐tuning of the progenitor binary (involving either a somewhat extreme mass ratio and/or mostly aligned spins). Our analysis shows that stellar evolution alone can naturally explain the primary BH in GW231123 without requiring a hierarchical-merger origin.
Outlook
@@
We performed the first investigation of stars above the upper edge of the BHMG that includes both the effect of stellar rotation and variations in the critical $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ nuclear-reaction rate. The main conclusion of this paper is that stellar evolution alone can explain the primary mass and spin of GW231123, suggesting that this may be the first detection of a BH originating from a photodisintegrated star.
Previous studies have shown that much of the information on population features often comes from a few highly informative events [46, 47, 48, 49, 50]. As was the case for GW190521 when analyzing data from the first three LIGO/Virgo/KAGRA observing runs [51], we anticipate that the exceptionally large masses and spins of GW231123 will play a major role in upcoming GW population constraints with O4 data. Accordingly, our conclusions that the large masses of GW231123 can be produced in combination with large spins through conventional stellar evolution provides a foundation for population‐level studies. We anticipate that it will be important for future GW population analyses to allow the stellar component to accommodate systems like GW231123, rather than imposing a hard cutoff at the BHMG. Fixing such a cutoff a priori could bias inferences toward exotic formation scenarios.
@@ Acknowledgements
We thank Monica Colpi for discussions. D.C. is supported by STFC Grant No. ST/T001011/1. J.S. is supported by NSF Grant No. 2207880. D.G. is supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, Italian-French University (UIF/UFI) Grant No. 2025-C3-386, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), MSCA Fellowship No. 101064542–StochRewind, MSCA Fellowship No. 101149270–ProtoBH, MUR Young Researchers Grant No. SOE2024-0000125, and the ICSC National Research Centre funded by NextGenerationEU. Our simulations were run on the University of Hawai i’s high-performance supercomputer KOA. The technical support and advanced computing resources from University of Hawai i Information Technology Services – Cyberinfrastructure, funded in part by the NSF MRI award No. 1920304, are gratefully acknowledged.

### v2

@@ 4.1 Impact of rotation on the black hole mass gap
The bottom panel of Fig. 1 presents these same models in the BH spin-mass plane. Stars with a helium core mass of $160{\rm M}_{\odot}$ cannot form BHs with a spin above a certain threshold, which depends on the $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ reaction rate. In particular, BHs with spins and masses compatible with the $90\%$ confidence interval of the primary in GW231123 cannot be formed if the $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ reaction rate is smaller than its measured median value.
@@
As anticipated above, our results demonstrate that the photodisintegration boundary which marks the upper edge of the BHMG shifts to higher stellar masses in the presence of rapid rotation. As shown in Fig. 2, rotating stars reach lower core temperatures $T_{c}$ before the onset of explosive oxygen burning leads to rapid expansion and a temperature drop. Thus, these stars avoid reaching the conditions required for photodisintegration-induced collapse, and instead undergo PISN. For the median $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate, we find that this occurs at $\Omega_{\rm PISN}\sim 0.4\Omega_{\rm crit}$.
@@
Figure 2: Central temperature $T_{c}$ versus time since core helium depletion for the simulations described in text, with $M_{\rm in}=160{\rm M}_{\odot}$ for the median $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate. Non-rotating models reach higher core densities and temperatures, crossing into the regime where photodisintegration reactions lead to gravitational collapse, here assumed to be $T_{c}=9\times 10^{9}\rm\,K$. Rotating models, by contrast, terminate at lower $T_{c}$, thereby avoiding collapse and instead undergoing a PISN.
4.2 Implications for GW231123
@@
Our simulations imply that it is plausible to interpret GW231123’s primary component ($m_{1}\simeq 137{\rm M}_{\odot}$) as having formed via photodisintegration instability collapse, even with a rapidly rotating core in the absence of efficient angular momentum transport in the star. A similar claim was previously made (Fishbach and Holz, 2020) for the primary BH in GW190521, the former high‐mass record holder among GW events, although that interpretation depends strongly on the adopted prior distribution when analyzing the underlying GW data.
@@
In our models, the $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ reaction rate needed to form GW231123’s primary is both broad and realistic, ranging from the median to rates $3\sigma$ stronger. The upper end is compatible with previous hints that higher rates are needed to explain GW observations e.g., the $35{\rm M}_{\odot}$ peak in the BH mass function (Croon and Sakstein, 2025).
While our analysis has focused on the primary component of GW231123, it is worth briefly commenting on the secondary BH. The inferred mass posterior for the secondary is compatible with formation below the gap — the lower edge of the 90% credibility interval is $m_{2}\sim 50\,M_{\odot}$. A natural possibility is therefore that GW231123 straddles the mass gap. This interpretation is supported from a GW data analysis perspective: in a massive event like GW231123, the total mass $m_{1}+m_{2}$ is measured more accurately than either individual mass, which are consequently anti-correlated. If $m_{1}$ lies toward the upper end of its uncertainty range (placing it above the gap), then $m_{2}$ will lie toward the lower end of its range (and thus below the gap), a mechanism previously suggested to explain GW190521 by Fishbach and Holz (2020). As shown in this paper, this picture is further supported by the fact that rotation shifts the lower edge of the mass gap to higher masses. The secondary BH does not pose additional challenges to stellar-evolution models, although a dedicated evolutionary study of this object would certainly be worth pursuing.
5 Outlook
@@
We performed the first investigation of stars above the upper edge of the BHMG that includes both the effect of stellar rotation and variations in the critical $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ nuclear-reaction rate. The main conclusion of this paper is that in the regime of weak angular momentum transport motivated by the spins of GW231123, it is possible that the primary component of GW231123 formed from direct stellar collapse.
An important assumption of this analysis is that we do not include the ST dynamo or other mechanisms that enforce strong angular momentum coupling. Efficient coupling would drive the core toward uniform rotation and produce nearly non-spinning black holes (e.g. Marchant and Moriya 2020; Marchant et al. 2024), which would be inconsistent with the high spin inferred for GW231123 under current interpretations. Our results should therefore be understood as an exploration of the BHMG in the presence of rapidly rotating stellar cores, motivated by proposed scenarios in which massive stars can retain or regain substantial angular momentum.
Previous studies have shown that much of the information on population features often comes from a few highly informative events (Gaebel et al., 2019; Essick et al., 2022; Moore and Gerosa, 2021; Baxter et al., 2021; Mancarella and Gerosa, 2025). As was the case for GW190521 when analyzing data from the first three LIGO/Virgo/KAGRA observing runs (Abbott and others, 2023), we anticipate that the exceptionally large masses and spins of GW231123 will play a major role in upcoming GW population constraints with O4 data. Accordingly, our conclusions that the large masses of GW231123 can be produced in combination with large spins through stellar evolution in the weak-coupling regime provides a foundation for population‐level studies. We anticipate that it will be important for future GW population analyses to allow the stellar component to accommodate systems like GW231123, rather than imposing a hard cutoff at the BHMG. Fixing such a cutoff a priori could bias inferences toward exotic formation scenarios.
@@ Acknowledgements
We thank Monica Colpi for discussions. D.C. is supported by STFC Grant No. ST/T001011/1. J.S. is supported by NSF Grant No. 2207880. D.G. is supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, Italian-French University (UIF/UFI) Grant No. 2025-C3-386, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), MSCA Fellowship No. 101064542–StochRewind, MSCA Fellowship No. 101149270–ProtoBH, MUR Young Researchers Grant No. SOE2024-0000125, and the ICSC National Research Centre funded by NextGenerationEU. J.S. thanks the IPPP for hospitality and acknowledges an IPPP DIVA fellowship to support the visit. J.S. and D.G. thank the organizers of the International Congress of Basic Science (Beijing, 2025). Our simulations were run on the University of Hawai i’s high-performance supercomputer KOA. The technical support and advanced computing resources from University of Hawai i Information Technology Services – Cyberinfrastructure, funded in part by the NSF MRI award No. 1920304, are gratefully acknowledged.
Data Availability
Our code can be found in the paper’s reproduction package at the following URL: https://zenodo.org/records/16898502. Software used includes: MESA version 15140, MESASDK version 20210401, GFORTRAN GCC version 9.2.0, GFORTRAN GCC version 9.2.0, Jupyter Notebook version 6.4.12, Python version 3.8.5, Mathematica via Wolfram version 14.2.
Appendix A Sensitivity to collapse description
In Fig. 3 we examine the dependence of the final BH mass and dimensionless spin parameter $\chi$ on the initial rotation rate $\Omega/\Omega_{\rm crit}$ for two different choices of the prompt‐collapse core mass, $M_{\rm cutoff}=3\,M_{\odot}$ and $5\,M_{\odot}$.  The left panel shows that across the full range $0.10\leq\Omega/\Omega_{\rm crit}\leq 0.40$, the resulting BH mass varies by at most $\Delta M\lesssim 0.4\,M_{\odot}$, while the right panel demonstrates that the spin parameter changes by less than $\Delta\chi\lesssim 0.02$.  These small offsets confirm that our assumption of the mass of the inner core collapsing promptly has a negligible impact on the median‐rate outcomes.
Figure 3: Final black‐hole mass (top) and dimensionless spin parameter $\chi$ (bottom) as functions of the initial rotation rate $\Omega/\Omega_{\rm crit}$ for two choices of the prompt‐collapse core mass threshold, $M_{\rm cutoff}=3\,M_{\odot}$ and $5\,M_{\odot}$